OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = -14 x = -17

OpenStudy (yttrium):

Is it, \[\frac{ x+20 }{ 2 } = 3x\]

OpenStudy (anonymous):

Connie has saved up $15 to purchase a new CD from the local store. The sales tax in her county is 5% of the sticker price. Write an equation and solve it to determine the value of the highest priced CD Connie can purchase with her $15, including the sales tax. Round your answer to the nearest penny. x + 0.05x = 15; x = $14.29 x - 0.05x = 15; x = $15.79 0.5x = 15; x = $30 5x = 15; x = $3.00

OpenStudy (yttrium):

I think it's the 1st equation

OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960